OpenStudy (anonymous):

If the perimeter of a square is one hundred forty-four miles, what is its area? twelve square miles forty-eight square miles one hundred forty-four square miles one thousand two hundred ninety-six square miles

OpenStudy (anonymous):

P = 4s 144 = 4s -- divide both sides by 4 36 = s A = s^2 A = 36^2 A = 1296

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

no problem :)